\begin{thebibliography}{99}

\bibitem{Arnold:89} V. I. Arnold,
\emph{Mathematical methods of classical mechanics.} Translated by K. Vogtman and A. Weinstein. (English)
Graduate Texts in Mathematics. \textbf{60}. New York - Heidelberg - Berlin: Springer-Verlag, (1989).

\bibitem{Berger:03} M. Berger,
A panoramic view of Riemannian geometry. 
Berlin: Springer, (2003).


\bibitem{BolMa/So:99} V. Boltyanski, H. Martini, V. Soltan, Geometric Methods and Optimization Problems, Kluwer, Dordrecht/Boston/London (1999).

\bibitem{Fermat:91} P. de. Fermat, Oeuvres de Fermat, vol.~1,Gauthiers-Villars, Paris (1891).



\bibitem{Kup/Mar:97} Y.S. Kupitz and H. Martini, \emph{Geometric aspects of the generalized Fermat-Torricelli problem},
Bolyai Society Mathematical Studies.\textbf{6} (1997) , 55-127.


\bibitem{Lagrange:}J.L.Lagrange,
Mécanique analytique. Tome 1, 2. Édition complète réunissant les notes de la troisième édition revue, corrigée et annotée par Joseph Bertrand et de la quatrième édition publiée sous la direction de Gaston Darboux. (French) 
Paris: Librairie Scientifique et Technique Albert Blanchard, (1965). 

\bibitem{Lindelof:67}L.L.Lindel$\ddot{o}$f, \emph{Sur les maxima et minima d’ une fonction des rayons vecteurs menés d’un point mobile à plusieurs centres fixes.} Acta Soc. Sc. Fenn. \textbf{8} (1867), 191-203.


\bibitem{Mach:19} E. Mach, The science of mechanics; a critical and historical account of its development, Chicago/London The Open Court Publishing (1919).
\bibitem{Sturm:84} R. Sturm, Ueber den Punkt kleinster Entfernungsumme von gegebenen Punkten.
J. Reine Angew. Math. XCVII (1884), 49-62.


%------------------------------------------------

%---------------



%-----------------------------------------------------------


%----------------------------------------------------------------------------

\end{thebibliography}

\documentclass[reqno]{amsart}
\usepackage{amsmath}
\usepackage[dvips]{graphicx}
\usepackage{amsfonts}
\usepackage{amssymb}
\usepackage{latexsym}
\graphicspath{{Img/}}
\newtheorem{theorem}{Theorem}
\theoremstyle{plain}

\newtheorem{corollary}{Corollary}

\newtheorem{lemma}{Lemma}

\newtheorem{problem}{Problem}

\numberwithin{equation}{section}

\begin{document}

\title[D'Alembert-Lagrange principle yields a weighted balancing condition]{D'Alembert-Lagrange principle for point masses yields a system of weighted balancing unit vectors  in the three dimensional Euclidean Space}
\author{Anastasios N. Zachos}
\address{University of Patras, Department of Mathematics, GR-26500 Rion, Greece}
\email{azachos@gmail.com}
\keywords{D'Alembert-Lagrange principle, point masses, weighted Fermat problem, weighted Fermat-Torricelli point, weighted balancing condition} \subjclass[2010]{Primary 70S05;37J06 Secondary 51M16}

\dedicatory{To Lagrange's superposition of equilibria}

\begin{abstract}
In this paper, we prove that D'Alembert-Lagrange principle for point masses using Lagrange-Mach's mechanical construction yields a weighted balancing condition of unit vectors in $\mathbb{E}^{3}.$
\end{abstract}\maketitle

\section{Introduction}

We start by giving D'Alembert-Lagrange principle for a system of points $x_{i}$ with masses $m_{i},$ which move freely in $\mathbb{E}^{3}$  in the configuration manifold $\mathbb{R}^{3n}$ for $i=1,2,\ldots,n$ (\cite[Section~21,pp.~91-97]{Arnold:89},\cite{Lagrange:}).

We note that tangent vectors to the configuration manifold are called virtual variations and the constraint forces are defined by $\textbf{R}_{i}:=m_{i}\ddot{x}_{i}+\frac{\partial U}{\partial x_{i}}.$

We denote by $TM_{x}$ the tangent space of the surface $M$ in the three dimensional space $\{x\}.$

The D'Alembert-Lagrange principle states that:

The sum of the works of the constrained forces on any virtual variation $\xi_{i}\in TM_{x}$ is zero or
$\sum_{i=1}^{n}(m_{i}\ddot{x}_{i}+\frac{\partial U}{\partial x_{i}},\xi)=0.$

We recall the weighted Fermat Problem for $m$ non-collinear and non-coplanar points $A_{i}(x_{i},y_{i},z_{i})$  in $\mathbb{E}^{3}$ (\cite{Sturm:84},\cite{Lindelof:67},\cite{Kup/Mar:97} and \cite{BolMa/So:99}), which was first states in the 17th century by Fermat for three unweighted points (\cite{Fermat:91}):

\begin{problem}[The weighted Fermat problem in $\mathbb{R}^{3}$]\label{WFN}
Find $A_{0}(x,y,z)$ in $\mathbb{R}^{3},$ such that:
\begin{equation}\label{objectivewfrn}
f(\{A_{0}\})=\sum_{i=1}^{m}b_{i}\sqrt{(x-x_{i})^2+(y-y_{i})^2+(z-z_{i})^2}\to min.
\end{equation}
\end{problem}

We need the following fundamental lemma, which deals with the characterization of solutions for the weighted Fermat problem (\cite{Kup/Mar:97},\cite{BolMa/So:99}). The unweighted case was first proved by Lindelof and Sturm in \cite{Lindelof:67} and \cite{Sturm:84}.



\begin{lemma}{\cite[Theorem~18.37,pp.~250]{BolMa/So:99},\cite{Kup/Mar:97}}\label{characteizationFTsolution}
(I) The weighted Fermat point $A_{0}$
exists and is unique.

(II) If for each point  $A_{i}\in \{A_{1},A_{2},...,A_{m}\}$
\[ \|\sum_{j=1,j\ne i}^{m}b_{j}u(A_{j},A_{i})\|>b_i, \]
for $i,j=1,2,\ldots,m,$ then

(a) the weighted Fermat (Fermat-Torricelli) point $A_{0}$ does not belong
to\\ $\{A_{1},A_{2},...,A_{m}\}$

(b) The following weighted balancing condition of unit vectors $u(A_{0},A_{i})$ occurs:

\[ \sum_{i=1}^{m}b_{i}u(A_{0},A_{i})=0\]
(weighted Fermat-Torricelli tree solution).
\end{lemma}


In this paper, we prove that D'Alembert-Lagrange principle for point masses having only one point that moves freely in $E^{3}$ implies a weighted balancing condition of unit vectors in the spirit of weighted Fermat-Torricelli tree solution in $\mathbb{E}^{3}.$

\section{The equivalence of D'Alembert-Lagrange principle and Fermat-Torricelli's weighted balancing condition in two dimensions}
In this section, we generalize Lagrange-Mach's mechanical construction for $m$ unequal weights in the Eucidean plane $\mathbb{E}^{2}$ and on a smooth surface $M$ in $\mathbb{E}^{3}$ and we prove the equivalence of D'Alembert-Lagrange principle and Fermat-Torricelli's weighted balancing condition.

Let $\{A_{1},A_{2},\ldots, A_{m}\}$ be $m$ fixed points and $A_{0}$ be a point in the Euclidean plane $\mathbb{E}^{2}$ or an horizontal plane and let pulleys be placed at these fixed points over which $m$ strings are passed loaded with weights $b_{1},b_{2},\ldots,b_{m},$ respectively and knotted at $A_{0}.$
We set $\alpha_{i0j}\equiv \angle A_{i}A_{0}A_{j}$  and the infinitesimal small distances $\delta \xi_{j}\equiv A_{0}A_{0^{\prime},j},$ such that
$A_{0^{\prime},j}\in [A_{0},A_{j}].$
We denote by $A_{0^{\prime},ij}$ the trace of the orthogonal projection of $A_{0^{\prime},j}$ to the line defined by $A_{0},A_{i}$ and by $A_{i^{\prime}}$ the symmetrical of $A_{i}$ with respect to $A_{0}.$

Furthermore, we set:

$c_{i,j}\equiv \begin{cases} -\cos\alpha_{j0i^{\prime}},& \text{if $A_{0^{\prime},ij}$
lies on the line segment [$A_{0},A_{i^{\prime}}$] },\\
\cos\alpha_{j0i},& \text{if $A_{0^{\prime},ij}$ lies on the line segment [$A_{0},A_{i}$]},
\end{cases}$

where $\alpha_{j0i^{\prime}}=\pi-\alpha_{j0i},$ for $i,i^{\prime},j=1,2,\ldots,m$ and $0<\alpha_{j0i^{\prime}},\alpha_{j0i}<\pi.$

\begin{theorem}\label{DalembertLagrangeMachR2}
D'Alembert Lagrange principle for Lagrange-Mach's mechanical construction of pulleys in $\mathbb{E}^{2}$ implies that:
\begin{equation}\label{sumweightedunitvectors}
\sum_{i=1}^{m}b_{i}u(A_{0},A_{i})=0.
\end{equation}
\end{theorem}

\begin{proof}

By displacing the point $A_{0}$ in the directions of $A_{0}A_{j}$ by the infinitesimal small distances $\delta \xi_{j},$ we obtain:

\[\sum_{j=1}^{m}\sum_{i=1}^{m}c_{i,j}b_{i}\delta \xi_{i}=0,\]

which gives

\begin{equation}\label{sumcosr2}
\sum_{i=1}^{m}b_{i} \cos\alpha_{j0i}=0,
\end{equation}

for every $j=1,2,\ldots,m.$

By taking a fixed index $k\in \{1,2,\ldots,m\}$ and the addition of angles $\angle A_{j0k}+\angle A_{k}A_{0}A_{i}$ in $\mathbb{E}^{2},$ we get:  
\[\sum_{i=1}^{m}b_{i}\cos(\alpha_{j0k}+\alpha_{k0i})=0\]

or

\[\sum_{i=1}^{m}b_{i}\cos\alpha_{j0k}\cos\alpha_{k0i}-\sum_{i=1}^{m}b_{i}\sin\alpha_{j0k}\sin\alpha_{k0i}=0,\]

which gives

\[\sum_{i=1}^{m}b_{i}\sin\alpha_{k0i}=0.\]

Thus, for $k=j,$ we get:

\begin{equation}\label{sumsinr2}
\sum_{i=1}^{m}b_{i}\sin\alpha_{j0i}=0
\end{equation}

By taking into account (\ref{sumcosr2}),(\ref{sumsinr2}) we derive (\ref{sumweightedunitvectors}).

\end{proof}

\begin{corollary}{Lagrange-Mach mechanical contsruction \cite[pp.~61-62]{Mach:19}}
For $b_{1}=b_{2}=b_{3}>0$ and $b_{4}=\ldots=b_{m}=0,$ we get
\[\sum_{i=1}^{3}u(A_{0},A_{i})=0.\]
\end{corollary}

Let $\{A_{1},A_{2},\ldots, A_{m}\}$ be $m$ fixed points and $A_{0}$ be a point on a smooth surface $M$ in $\mathbb{E}^{3}$  and let pulleys be placed at these fixed points over which $m$ strings are passed loaded with weights $b_{1},b_{2},\ldots,b_{m},$ respectively and knotted at $A_{0}.$ Furhtermore, the loose ends are passed through $m$ holes, which correspond to the fixed points $\{A_{1},A_{2},\ldots, A_{m}\}$ and are attached to the physical weights
$b_{1},b_{2},\ldots,b_{m},$ respectively.

We assume the existence of a small positive real number $I,$ such that the  injectivity radius of $M$ $inj(M):=I.$ Therefore, we may consider a bounded and connected region $D$ on $M$, which is subset of a disk $B_{X;I}$ with center $X$ and radius $R=I,$ such that $\{A_{0,}A_{1},A_{2},\ldots, A_{m}\}$ in $D.$
Hence, any geodesic arc that connects $\{A_{0,}A_{1},A_{2},\ldots, A_{m}\}$ with any point $Y$ in $D$ is a shortest arc (segment) (\cite[p.~42]{Berger:03}).
Thus, we define the exponential map at $A_{0}$ $exp_{A_{0}}:W_{A_{0}}\subset T_{A_{0}}(M)\to M$ from the vector $w$ starting at $A_{0}$ travelling along up to the point $ \gamma_{\frac{w}{||w||}}(||w||)$ having length $||w||$ (\cite[pp.~222-223]{Berger:03}).

\begin{theorem}\label{DalembertLagrangeMachM}
D'Alembert Lagrange principle for Lagrange-Mach's mechanical construction of pulleys on a smooth surface $M$ in $\mathbb{E}^{3}$ implies that:
\begin{equation}\label{sumweightedunitvectorssurface}
\sum_{i=1}^{m}b_{i}\frac{exp_{A_{0}}^{-1}(A_{0},A_{i})}{||exp_{A_{0}}^{-1}(A_{0},A_{i})||}=0.
\end{equation}
\end{theorem}

\begin{proof}
By applying Lagrange-Mach's mechanical construction of pulleys on a smooth surface $M$ in $\mathbb{E}^{3},$ we obtain equilibrium at $A_{0},$ such that
the $m$ strings are attached at the geodesic arcs $A_{i}A_{0}.$ By taking the inverse of the exponential mapping at $A_{0}$ $exp_{A_{0}}^{-1}(A_{i}),$ we may move to $T_{A_{0}}(M).$ Hence, by applying Theorem~\ref{DalembertLagrangeMachR2}, we obtain a system of weighted balancing unit vectors at $T_{A_{0}}(M),$ which gives (\ref{sumweightedunitvectorssurface}).   
\end{proof}

\section{The equivalence of D'Alembert-Lagrange and Fermat-Torricelli's weighted balancing condition in three dimensions}
In this section, we generalize Lagrange-Mach's mechanical construction for $m$ unequal weights in the three dimensional Eucidean Space $\mathbb{E}^{3}$ and we prove the equivalence of D'Alembert-Lagrange principle and Fermat-Torricelli's weighted balancing condition.

Let $\{A_{1},A_{2},\ldots, A_{m}\}$ be $m$ fixed points and $A_{0}$ be a point in $\mathbb{E}^{3}$ and let pulleys be placed at these fixed points over which $m$ strings are passed loaded with weights $b_{1},b_{2},\ldots,b_{m},$ respectively and knotted at $A_{0}.$
We set the infinitesimal small distances $\delta \xi_{j}\equiv A_{0}A_{0^{\prime},j},$ such that
$A_{0^{\prime},j}\in [A_{0},A_{j}].$

We denote by $A_{i^{\prime}}$ the symmetrical point of $A_{i}$ with respect to $A_{0},$ by $A_{k,i0j}$ the trace of the orthogonal projection of $A_{k}$ to the plane defined by $\triangle A_{i}A_{0}A_{j},$ by $A_{k^{\prime},i0j}$ the symmetrical point of $A_{k,i0j}$ with respect to $A_{0},$ by $A_{0^{\prime\prime},(0k^{\prime},i0j)}$ the trace of the orthogonal projection of $A_{0^{\prime},j}$ to the line defined by the segment $A_{0}A_{k^{\prime},i0j}$ and by $A_{0^{\prime\prime\prime},0k^{\prime}}$ the trace of the orthogonal projection of $A_{0^{\prime\prime},(0k^{\prime},i0j)}$
to the line defined by the segment $A_{0}A_{k^{\prime}},$ for $i,j,k,i^{\prime},k^{\prime}=1,2,\ldots,m.$

We set:

$\alpha_{i,j0k}\equiv \angle A_{i}A_{0}A_{i,j0k},$

$\omega_{i,j0k}\equiv \angle A_{i,j0k}A_{0}A_{j},$

$c_{i,j0k}=\begin{cases} -\delta \xi_{k}\cos\omega_{i,j0k}\cos\alpha_{i,j0k},& \text{if $A_{0^{\prime\prime\prime},0i} \notin$ [$A_{0},A_{i}$] },\\
\delta \xi_{k}\cos\omega_{i,j0k}\cos\alpha_{i,j0k},& \text{if $A_{0^{\prime\prime\prime},0i}\in [A_{0},A_{i}]$}.
\end{cases}$


\begin{theorem}\label{DalembertLagrangeMachR3}
D'Alembert Lagrange principle for Lagrange-Mach's mechanical construction of pulleys in $\mathbb{E}^{3}$ implies that:
\begin{equation}\label{sumweightedunitvectorsr3}
\sum_{i=1}^{m}b_{i}u(A_{0},A_{i})=0.
\end{equation}
\end{theorem}

\begin{proof}

We consider the following system of unit vectors $u(A_{0},A_{i})$ in $\mathbb{E}^{3}:$

We express the unit vectors $\vec {u}(A_{0},A_{i})$ for $i=1,2,3,4,$ using spherical coordinates:

\begin{equation}\label{spherical1}
u(A_{0},A_{j})=(1,0,0),
\end{equation}

\begin{equation}\label{spherical2}
u(A_{0},A_{k})=(\cos\alpha_{j0k},\sin\alpha_{j0k},0),
\end{equation}

\begin{equation}\label{spherical3}
u((A_{0},A_{i})=(\cos a_{i,j0k} \cos\omega_{i,j0k},\cos a_{i,j0k} \sin\omega_{i,j0k},\sin a_{i,j0k} ),
\end{equation}

$\ldots$

\begin{equation}\label{spherical4}
u(A_{0},A_{m})=(\cos a_{m,j0k} \cos\omega_{m,j0k},\cos a_{m,j0k} \sin\omega_{m,j0k},\sin a_{m,j0k} ).
\end{equation}

for $i,j,k=1,2,\ldots,m, i\neq j\ne k.$

By displacing the point $A_{0}$ in the directions of $A_{0}A_{j}$ by the infinitesimal small distances $\delta \xi_{j},$ we obtain:

\[\sum_{k=1}^{m}\sum_{i=1}^{m}c_{i,j0k}b_{i}\delta \xi_{k}=0,\]

which gives

\begin{equation}\label{sumcosr3}
\sum_{k=1}^{m}\sum_{i=1}^{m}b_{i} \cos\alpha_{i,j0k}\cos\omega_{i,j0k}=0,
\end{equation}

or

\begin{equation}\label{sumcosr3bis}
\sum_{i=1}^{m}b_{i} \cos\alpha_{i,j0k}\cos\omega_{i,j0k}=0,
\end{equation}

for every $j=1,2,\ldots,m.$

By substituting the inner product 
\[u(A_{0},A_{j})\cdot u(A_{0},A_{k})=\cos\alpha_{j0k}, \]
in (\ref{sumcosr3}), we obtain:

\begin{equation}\label{sumcosanglesr3}
\sum_{k=1}^{m}b_{i}\cos\alpha_{j0k}=0.
\end{equation}

Without loss of generality, we fix the indices $j=1, k=2$ and we consider the particular system of unit vectors:

\begin{equation}\label{spherical1102}
u(A_{0},A_{1})=(1,0,0),
\end{equation}

\begin{equation}\label{spherical2102}
u(A_{0},A_{2})=(\cos\alpha_{102},\sin\alpha_{102},0),
\end{equation}

$\ldots$

\begin{equation}\label{spherical3102}
u((A_{0},A_{k})=(\cos a_{k,102} \cos\omega_{k,102},\cos a_{k,102} \sin\omega_{k,102},\sin a_{k,102} ).
\end{equation}

for $k=1,2,\ldots, m.$

By replacing $j=2$ and $j=3$ in (\ref{sumcosanglesr3}),

we get

\begin{equation}\label{sumcosanglesr320k}
\sum_{k=1}^{m}b_{i}\cos\alpha_{20k}=0.
\end{equation}

and

\begin{equation}\label{sumcosanglesr320k2}
\sum_{k=1}^{m}b_{i}\cos\alpha_{30k}=0.
\end{equation}

By substituting the inner product \[u(A_{0},A_{2})\cdot u(A_{0},A_{k}) =\cos\alpha_{102}\cos a_{k,102}\cos\omega_{k,102}+\sin\alpha_{102}\cos a_{k,102} \sin\omega_{k,102}\]
in (\ref{sumcosanglesr320k}) and taking into account (\ref{sumcosr3bis}), we derive:

\begin{equation}\label{sumcossin102r3}
\sum_{k=1}^{m}b_{i} \cos\alpha_{k,102}\sin\omega_{k,102}=0.
\end{equation}

By substituting the inner product 
\[u(A_{0},A_{3})\cdot u(A_{0},A_{k}) =\cos a_{3,102}\cos\omega_{3,102}\cos a_{k,102}\cos\omega_{k,102}+\]\[+\cos a_{3,102} \sin\omega_{3,102}\cos a_{k,102} \sin\omega_{k,102}+\sin\alpha_{3,102}\sin\alpha_{k,102}\]

in (\ref{sumcosanglesr320k2}) and taking into account (\ref{sumcosr3bis}), (\ref{sumcossin102r3}), we derive:

\begin{equation}\label{sumsin102r3}
\sum_{k=1}^{m}b_{i} \sin\alpha_{k,102}=0.
\end{equation}
Hence,  we obtain the following system of three equations (\ref{sumcosr3bis}),(\ref{sumcossin102r3}), (\ref{sumsin102r3}):

\[\sum_{k=1}^{m}b_{i} \cos\alpha_{k,102}\cos\omega_{k,102}=0,\]

\[\sum_{k=1}^{m}b_{i} \cos\alpha_{k,102}\sin\omega_{k,102}=0,\]

\[\sum_{k=1}^{m}b_{i} \sin\alpha_{k,102}=0,\]

which yields (\ref{sumweightedunitvectorsr3}).

\end{proof}

\end{document}